# Learning Concept-Driven Logical Rules for Interpretable and Generalizable Medical Image Classification

### v1

@@ 3.2.2 Model generalizability analysis:
To evaluate generalizability, we assess the out-of-domain (OOD) performance of the models on the unseen DDI dataset, using models trained on the source F17k dataset. As reported in Table 2, we can observe that concept models relying on soft concepts (probabilities or embeddings) exhibit relatively large performance drops, whereas methods employing binary concept values experience smaller declines. This suggests that binary concepts help mitigate concept leakage. Notably, although both CRL and hard-CBM utilize binary concepts, CRL achieves a diagnostic ACC of 73.46, outperforming the second-best method by 9.21 while exhibiting the smallest performance drop. This improvement can be attributed to the domain-invariant logical rules extracted by logical layers, which enhance robustness to distribution shifts.
@@ 3.2.3 Model interpretability analysis:
To illustrate that CRL could generate meaningful logical rules, we present the rules obtained from both skin disease diagnosis and WBC classification tasks, as illustrated in Fig. 3. For skin disease diagnosis, we observe that concepts Telangiectasia, Ulcer and Crust are associated with Malignant, aligning with established clinical knowledge [3]. In case of WBC classification, the rules indicate that Lymphocytes are characterized by a High NC ratio and a Round nucleus, while Neutrophils typically exhibit Pink granules and Dense chromatin, consistent with clinical observations [21]. The case studies demonstrate that CRL can effectively extract concept-based logical rules that are clinically meaningful.
@@ 4 Conclusion
This paper proposes CRL, a framework for interpretable medical image classification that mitigates concept leakage and unifies local and global explanations. By leveraging binary concepts and learnable logical layers, CRL effectively models concept correlations and extracts clinically meaningful decision rules. Experiments on two tasks demonstrate that CRL not only achieves competitive performance but also exhibits superior generalizability to unseen data.

### v2

@@ 3.2.2 Model generalizability analysis:
To evaluate generalizability, we assess the out-of-domain (OOD) performance of the models on the unseen DDI dataset, using models trained on the source F17k dataset. As reported in Table 2, we can observe that concept models relying on soft concepts (probabilities or embeddings) exhibit relatively large performance drops, whereas methods employing binary concept values experience smaller declines. This suggests that binary concepts could help mitigate concept leakage. Notably, although both CRL and hard-CBM utilize binary concepts, CRL achieves a diagnostic ACC of 73.46, outperforming the second-best method by 9.21 while exhibiting the smallest performance drop. This improvement can be attributed to the domain-invariant logical rules extracted by logical layers, which enhance robustness to distribution shifts.
@@ 3.2.3 Model interpretability analysis:
To illustrate that CRL could generate meaningful logical rules, we present the rules obtained from both skin disease diagnosis and WBC classification tasks, as illustrated in Fig. 3. For skin disease diagnosis, we observe that concepts Telangiectasia, Ulcer and Crust are associated with Malignant, aligning with established clinical knowledge [3]. In case of WBC classification, the rules indicate that Lymphocytes are characterized by a High NC ratio and a Round nucleus, while Neutrophils typically exhibit Pink granules and Dense chromatin, consistent with clinical observations [22]. The case studies demonstrate that CRL can effectively extract concept-based logical rules that are clinically meaningful, offering both local concept explanations and global rule explanations for the entire medical dataset.
@@ 4 Conclusion
This paper introduces CRL, a framework for interpretable medical image classification that mitigates concept leakage and unifies local and global interpretability. By employing binary concepts and learnable logical layers, CRL effectively models concept correlations and extracts clinically meaningful decision rules. Experiments on two medical image classification tasks demonstrate that CRL achieves competitive performance and exhibits superior generalizability to unseen data.